find the solution to each system by the substitution method. 3x-5y=7 x-4y= -7
Answer by rfer(16322)

x-4y=-7
x=4y-7
----------
3x-5y=7
3(4y-7)-5y=7
12y-21-5y=7
7y=28
y=28/7
y=4
-------------
3x-5(4)=7
3x-20=7
3x=27
x=27/3
x=9
(9,4)
----------------
x-4y=-7
9-16=-7
-7=-7